OpenStudy (anonymous):

Solve 2 – 3x = x – 10

OpenStudy (anonymous):

-4x=-12 x=3

OpenStudy (diyadiya):

1.Subtract x from both sides 2.Subtract 2 on both sides

OpenStudy (anonymous):

take x to left side and 2 to right and solve u will get 3

OpenStudy (diyadiya):

1.Subtract x from both sides 2-3x = x-10 2-4x =-10 2.Subtract 2 on both sides -4x =-10-2 -4x=-12 Now divide both sides by -4

OpenStudy (anonymous):

@claireh12 understood